How do you know which number is the dependent variable?
If you are given an equation, y is always the dependent variable. That's why you write y=something, to show that x changes and forces y to change too.
If you are given a word question, it becomes more complicated. You have to see which one is forced to change as the other changes, and which one is changing. For example, if the two variables are the speed of the car and the time it takes to stop completely, the speed is independent. You can figure this out by realizing that it makes no sense to know how long it takes to brake without knowing how fast you are going. Really, with word problems, it becomes a matter of common sense and experience.
Two mutually exclusive outcomes. You flip a coin, and only heads and tails are possible.
The person in knowledge is a better person as he knows more about theprinciples that govern the material, metaphysical and spiritual worlds andhence fully adheres to those principles n gets better results.
The person in passion may have the same knowledge but in thehaste of acting misses out on some elements of that action and gets a lesserresult.
For example if I plant an Alphanso mango seed and nurture it organically - ittakes 7 plus years before I get to enjoy the ripened and most tasty fruit. Andafter that every alternate year!
If I have the passionate nature, I will feverishly add chemicals etc n get toenjoy the ripened fruit in 4 odd years and every year there after- but thenatural taste is lost- its not so tasty!
The person in goodness/knowledge is calm n collected in his actions anddoes them in proper sequence and in knowledge and thereby obtains full resultswhere as the person in passion is in a hurry and will definitely miss out on afew sequential actions thereby losing out on achieving the correct result.
So to get best results one must act in full knowledge- calmly.
There is one rider though about the later type of person- that person could bein ignorance!
The external manisfestation of an ignorant person and a person in fullknowledge appears to be the same- the results are totally different.
For example the person sweeping the floor in goodness or in ignorance appear Tobe doing the same thing- sweeping the floor slowly but the person in knowledgeis sweeping away the dirt in one direction till it is collected into a dust panwhere as the person in ignorance is sweeping slowly but pushing some dirt backinto the swept area and doing the sweeping action again and again and missingout on some dirt- the result one sees in the dust collected and what remains on the floor!
I prefer to be the person who takes time to live a slower paced life but inknowledge- I will definitely achieve fulfillment and greater self satisfaction.
That is an intelligent way of living.
kanai
How many 5 digit numbers can you make with the digits 1-8?
Are you allowed to use a digit more than once? I will assume not:
There are 8 choices for the first digit. One digit has been used, so there are only 7 choices for the 2nd digit. ... 6 for the 3rd, etc.
The total number of numbers is 8x7x6x5x4 = 56x30x4 = 56x120 = 6720.
If you are allowed to repeat digits, then the answer is 85 = 32,768.
What is the difference between the smallest 5 digit odd number and the greatest 4 digit even number?
The smallest 5 digit odd number is: 10001
The greatest 4 digit even number is: 9998
Subtracting 9998 from 10001 gives you 3.
Disclaimer: This is not an easy answer because a lot depends on the sample size, distribution of population, etc. But for the sake of this I will assume that we are talking about a normally-distributed population and a sufficiently large sample size.
You want to test the hypothesis that smokers need less sleep than the general public. You are told that the general public needs an average of 7.7 hours of sleep. So:
Avgpop = 7.7
The hypothesis that you are testing (H1) is that smokers need less sleep:
Avgsmoker < Avgpopulation or Avgsmoker < 7.7 hours
Your null hypothesis (Ho) would be that there is no difference between the two groups OR that smokers need MORE sleep than the general population:
Avgpopulation >= Avgsmoker or Avgsmoker <= 7.7 hours
You test that hypothesis by taking a sample of X smokers and comparing their average number of sleep to that of your population. You then test to see whether any differences you observe would likely have been found due to chance or because the sample is different from the population. The more smokers that you randomly sample, the more likely it will be that a given difference will be due to smokers need a different amount of sleep than the general public.
Lets say that you sampled 200 smokers and their average night of sleep was 7.55 hours with a standard deviation of 1.0. Now we can calculate the test statistic. Because we are assuming a normal population and the sample size is greater than 30, we can use a one-sample z-test (it is one sample because we are taking one sample and comparing it to a known mean).
Z = (Sample Mean - Population Mean)/(Standard Deviation of Sample/Square Root of Sample Size)
Z = (7.55 - 7.7)/(1.0/Sqrt(200))
Z = -.15/.0707
Z = -2.12
We now compare our test statistic to a critical value table. (one for the z-statistic can be found here: http://math2.org/math/stat/distributions/z-dist.htm
That Z-statistic corresponds to a p-value of .017. This means that the probability of randomly drawing a sample with this observed average assuming the null hypothesis is true is .017 or 1.7%. At a confidence level of 98.3% and below you would reject the null hypothesis. THIS DOES NOT MEAN WITH CERTAINTY THAT SMOKERS NEED LESS SLEEP THAN THE GENERAL PUBLIC, however the data would suggest that the hypothesis that smokers need less sleep is true (in this case).
What are the advantages of using charts?
It is advantageous to use charts because it makes it easier to compare two variables over a period of time. It also becomes easier to explain or to present certain ideas.
Median is the middle value in a set of numbers. The median of 278 is 278.
What are possible patient-centered outcomes in Describe Verbalize List and Demonstrate?
write possible patient-centered outcomes using following terms: verbalize, list and demonstrate
Probability of drawing a red marble = 4/16 = 1/4
Probability of drawing not a red marble = 1 - 1/4 = 3/4
What is the probability of rolling 3 dice and getting a 2 a 4 and a 6?
The probability of getting any one number on a single die is 1/6. The three dice behave independently; hence, these the probabilities of these events can be multiplied.
(1/6)3
For the set 3 4 5 8 x the mean median and mode all have the same value what is the value of x?
For the set 3 4 5 8 5, the mean median and mode all have the same value.
x X over 2
= 2
Error has different meanings in different contexts. In some cases it is a synonym for "mistake". In this sense "Human Error" would be careless accidents that could be avoided with care. In Scientific terms, however, "error" simply refers to the difference between an observed value and it's theoretical "expected" value, regardless of whether this difference relates to a "mistake" by the experimenter, or is the result of some uncontrolled variable in the experiment.
When you graph the motion of an object you put on the horizontal axis and on the vertical axis?
On the horizontal axis you would probably plot the time. On the vertical axis you could plot displacement, velocity or acceleration.